Kirk Van​ Houten, who has been married for 23 ​years, would like to buy his wife an expensive diamond ring with a platinum setting on their 30​-year wedding anniversary. Assume that the cost of the ring will be ​$11,175 in 7 years. Kirk currently has ​$4,750 to invest. What annual rate of return must Kirk earn on his investment to accumulate enough money to pay for the​ ring?
Answer #1
$11,175 is the FV and $4,750 is the PV, 7
being n.
So we have,
11175 = 4750*(1+r)^7
Solving for n
n = (11175/4750)^(1/7)-1 = 13.00%
CHECK:
4750*1.13^7 = $      11,175